# Emergent Interpretable Symbols and Content-Style Disentanglement via Variance-Invariance Constraints

### v1

@@ 4.3 Learning Digits and Colors from Written Digit Strings
Dataset: We synthesize an image dataset of written digit strings on light backgrounds using all 10 digits and 8 different ink colors. The order of digits is random. All images are diversified with Gaussian noises, random blur, and foreground and background color jitters. More details about the dataset can be found in the appendix.
Figure 5: t-SNE visualization of the learned digit (content) and color (style) representations. The first row shows content representations before quantization, and the second row’s visualizations use the same colors as the digits are written.
Figure 6: Confusion matrices of learned codebooks for digits. The horizontal axes show digit labels from “0” to “9”, and the vertical axes show codebook atoms aligned with digit labels.
Experiment Results: The quantitive results can be seen in Table 3. Similar to the music task, V3 performs the best considering both representation disentanglement and codebook interpretability. Even though some unsupervised baselines also achieve a high codebook accuracy at large $K$ values, they fall behind in terms of latent space retrieval, especially in style. We visualize the latent representations in Figure 5, and compare the confusion matrices in Figure 6 for intuitive understanding.
Ablation Study: Table 4 shows the ablation results of V3 on the digit string dataset. The inferior performance of $\nu_{k}=\mathrm{SD}$ supports our assumption of its weakness in capturing multi-peak high-dimensional distributions. We observe that V3 trained without $\mathcal{L}_{\mathrm{fragment}}$ achieves the best codebook interpretability. Referring to the ablation in section 4.2, it is conceivable that a three-loss target can sometimes also satisfy the four constraints as defined in V3, and the optimal value for hyperparameters r and $\beta$ may differ by datasets, which marks a possible direction for future improvements.
@@ 5 Limitation
We have identified several limitations in our V3 method that necessitate further investigation. First, while V3 achieves good disentanglement and symbolic interpretability, it is not flawless — samples of different contents (say images of “8" and “9") may sometimes be projected into the same latent code. Inspired by human learning, which effectively integrates both mode-1 and mode-2 cognitive processes, we aim to enhance V3 by incorporating certain feedback or reinforcement. This adaptation could also facilitate the application of V3 to more complex domains such as general image or video. Additionally, V3 is currently optimized to disentangle content and style from data samples that include defined fragments. Extending this capability to unsegmented data, such as continuous audio, represents a significant area for future development. Furthermore, V3 assumes that content elements do not overlap, which does not hold in cases of polyphonic music or mixed audio. Addressing this challenge will require a more sophisticated approach that considers the hierarchical nature of content.

### v2

@@ 4.1 Datasets
Written Phone Numbers Dataset (PhoneNums): We synthesize an image dataset of written digit strings on light backgrounds using 8 different ink colors, mimicking a scenario of handwritten phone numbers. The order of digits is random. All images are diversified with noises, blur, and foreground and background color jitters. Models should learn digits and colors as content and style.
Monophonic Instrument Notes Dataset (InsNotes): We synthesize a dataset consisting of 16kHz monophonic music audio of 12 different instruments playing 12 different pitches in an octave. Every pitch is played for one second with a random velocity and amplitude envelope. The audio files are then normalized and processed to magnitude spectrograms. Models should learn pitches and timbres as content and style, respectively.
Street View House Numbers (SVHN) (Netzer et al., 2011): We select all images with more than one digit from the SVHN dataset. We crop the images to the bounding boxes of the digits and resize them to 32$\times$48. Models should learn digits as content, their fonts, texture and colors as style. Note that the styles can be seen as from a continuous space, and the fonts in SVHN are very diverse.
Sprites with Actions Dataset (Sprites) (ope, ; Yingzhen and Mandt, 2018): The Sprites dataset contains animated cartoon characters with random appearances. We use a modified version of the dataset taking video sequences of characters performing 9 different actions in random order. Models should learn the actions as content and appearances as style. Note that the styles can be seen as from a continuous space.
Librispeech Clean 100 Hours (Libri100) (Panayotov et al., 2015): Librispeech is a large-scale multi-speaker corpus of read English speech in various accents. We use the “clean” pool of the Librispeech dataset, where we select the 100-hour subset for training. We align the audio to 39 phonemes (24 consonants and 15 vowels) using ground truth transcriptions with the Montreal Forced Aligner (McAuliffe et al., 2017), then extract 80-dimensional log-mel spectrograms and resize them to a length of 64 frames. Models should learn phonemes as content and speakers’ voice as style. The styles are considered as from a continuous space as the speakers’ voices are very diverse.
4.2 Results of Content-Style Disentanglement
Table 2: Evaluation of pitch and timbre disentanglement on InsNotes using latent retrieval. Values are reported in percentage.
Method $K$ Content Style
PR-AUC Best F1 PR-AUC Best F1
${\bm{z}}^{\mathrm{c}}\uparrow$ ${\bm{z}}^{\mathrm{s}}\downarrow$ ${\bm{z}}^{\mathrm{c}}\uparrow$ ${\bm{z}}^{\mathrm{s}}\downarrow$ ${\bm{z}}^{\mathrm{c}}\downarrow$ ${\bm{z}}^{\mathrm{s}}\uparrow$ ${\bm{z}}^{\mathrm{c}}\downarrow$ ${\bm{z}}^{\mathrm{s}}\uparrow$
V3 $12$ 89.9 8.9 90.1 15.1 9.3 87.5 15.0 88.0
$24$ 76.2 8.7 80.0 14.2 12.8 68.9 20.3 70.0
$48$ 72.2 8.4 74.4 14.2 12.3 72.2 22.0 71.5
MINE-based $12$ 56.4 7.61 62.0 14.2 10.3 61.4 16.9 63.7
$24$ 50.5 8.5 59.1 14.9 14.7 53.4 19.5 51.4
$48$ 44.6 10.2 54.0 16.5 13.8 52.1 18.3 49.7
Cycle loss $12$ 49.7 8.7 57.9 15.2 10.7 12.7 18.2 19.0
$24$ 47.0 8.7 54.5 15.2 14.2 18.9 19.4 23.1
$48$ 42.4 8.0 49.4 14.5 16.2 20.0 22.4 24.4
$\beta$-VAE - 18.1 20.8 12.2 19.0
EC2-VAE (c) - 83.2 8.0 86.2 14.2 10.7 60.0 16.9 62.8
EC2-VAE (c & s) - 90.4 7.9 90.4 14.2 11.1 90.5 18.0 90.4
On PhoneNums and InsNotes where concrete style labels are available, we evaluate the models’ content-style disentanglement ability by conducting a retrieval experiment to examine the nearest neighbors of every input ${\bm{z}}^{\mathrm{c}}$ and ${\bm{z}}^{\mathrm{s}}$ using ground truth content and style labels, evaluated by the area under the precision-recall curve (PR-AUC) and the best F1 score. We experiment with different codebook sizes $K$ to allow different levels of vocabulary redundancy. The results are shown in Table 1 and Table 2. We see that V3 outperforms unsupervised baselines on both datasets, and the performance is consistent across different codebook sizes $K$. V3 also outperforms EC2-VAE (c) in the style retrieval task, which indicates that V3 learns better-disentangled style representations containing less content information. Visualizations of content and style latent representations learned by V3 also show clearer grouping compared to baselines, for which we refer readers to Appendix C.1.
Table 3: Linear probing accuracies (in %) for content (digit) classification on SVHN. Method $K$ ${\bm{z}}^{\mathrm{c}}\uparrow$ ${\bm{z}}^{\mathrm{s}}\downarrow$ V3 $20$ 40.6 18.5 MINE-based $20$ 36.0 20.8 Cycle loss $20$ 16.8 21.2 $\beta$-VAE - 21.8 Raw input - 21.4 EC2-VAE (c) - 97.0 21.2    Table 4: Linear probing accuracies (in %) for content (action) classification on Sprites. Method $K$ ${\bm{z}}^{\mathrm{c}}\uparrow$ ${\bm{z}}^{\mathrm{s}}\downarrow$ V3 $18$ 88.2 20.2 MINE-based $18$ 79.1 22.2 Cycle loss $18$ 86.4 39.7 $\beta$-VAE - 33.2 Raw input - 99.0 EC2-VAE (c) - 99.8 15.7
On SVHN and Sprites where there are no style labels, we evaluate the models’ disentanglement ability by linear probing on the learned representations to predict content labels. We also compare with a linear classifier on raw input features. The classifier layer is trained for one epoch before evaluated on the test set. On both datasets we allow a 100% content vocabulary redundancy, resulting in $K=20$ for SVHN and $K=18$ for Sprites. The resulting accuracies are shown in Table 4 and Table 4. V3 outperforms unsupervised baselines on both datasets, only trailing behind the weakly supervised EC2-VAE (c) as the latter’s ${\bm{z}}^{\mathrm{c}}$ space is optimized for discriminative task.
Table 5: Linear probing accuracies (in %) for content (phoneme) classification on Libri100. Method $K$ ${\bm{z}}^{\mathrm{c}}\uparrow$ ${\bm{z}}^{\mathrm{s}}\downarrow$ V3 $80$ 52.1 40.4 MINE-based $80$ 28.6 51.6 Cycle loss $80$ 16.1 50.5 $\beta$-VAE - 11.0 Raw input - 31.8 EC2-VAE (c) - 78.1 18.2    Table 6: Speaker verification equal error rates (in %) with average embedding on Libri100. Method $K$ ${\bm{z}}^{\mathrm{c}}\uparrow$ ${\bm{z}}^{\mathrm{s}}\downarrow$ V3 $80$ 49.5 42.5 MINE-based $80$ 49.9 45.2 Cycle loss $80$ 49.8 45.9 $\beta$-VAE - 50.0 Raw input - 47.1 EC2-VAE (c) - 49.2 49.6
On Libri100, we evaluate the disentanglement ability by linear probing on the learned representations to predict content labels, as well as conducting a vanilla speaker verification experiment using the average embeddings of fragments in every utterance. We also allow a content vocabulary redundancy of about 100% ($K=80$). The results are shown in Table 6 and Table 6. The content and style embeddings learned by V3 shows better performance on their respective tasks and lower performance on the other task, indicating that V3 learns better disentangled representations.
4.3 Content Classification on Out-of-distribution Styles
We further evaluate the generalization ability of V3 on PhoneNums and InsNotes by testing the models’ content classification performance on a special test set with only unseen styles, provided with few-shot examples. We focus on comparing V3 with the weakly supervised method EC2-VAE (c) and a pure CNN classifier to evaluate the generalization ability introduced by latent disentanglement. In the $n$-shot settings, models are presented with $n$ samples of each content and new style combination. All models are continuously trained on new samples until performance stops improving. For V3, we choose the V3 versions with no codebook redundancy for comparison ($K=10$ for PhoneNums, $K=12$ for InsNotes) as they show a one-to-one mapping from codebook entries to content labels (see Section 4.4 and Appendix C.2 for details). We first align the learned codebook entries to ground truth content labels, and obtain classification results by the encoded content representations ${\bm{z}}^{\mathrm{c}}$. For EC2-VAE, we try two different continuous training strategies: 1) using pseudo content labels from its own predictions for self-boosting, as well as training the reconstruction loss, and 2) only optimize the reconstruction loss. Additionally, we compare with EC2-VAE and the CNN classifier provided with labels in continuous training. The results are shown in Table 7. Although V3 might fall behind supervised methods in the 0-shot setting, it comes to the lead in few-shot settings on both datasets as the number of extra samples increases. This indicates that V3 can learn by itself to make sense of unseen styles with only a few examples, an ability that emerges from learning representations and disentangled interpretable factors.
Table 7: Content classification accuracies (in %) on data with OOD styles.
Pretraining Continuous Training PhoneNums InsNotes
Method Supervision Supervision Self-boost 0-shot 1-shot 5-shot 10-shot 0-shot 1-shot 5-shot 10-shot
V3 No No No 57.8 91.3 97.1 99.0 90.5 97.6 97.8 99.2
EC2-VAE (c) Yes No No 84.2 92.1 92.2 92.7 87.1 87.2 89.4 91.2
EC2-VAE (c) Yes No Yes 84.2 91.8 92.1 92.4 87.1 94.6 95.0 95.1
CNN Classifier Yes No No 59.5 59.5 59.5 59.5 92.6 92.6 92.6 92.6
CNN Classifier Yes No Yes 59.5 80.2 82.2 82.7 92.6 87.6 85.9 85.3
EC2-VAE (c) Yes Yes No 84.2 94.6 98.8 99.2 87.1 97.7 98.9 99.8
CNN Classifier Yes Yes No 59.5 81.2 82.4 83.5 92.6 91.9 91.3 89.1
4.4 Results of Symbolic Content Interpretability
We notice that interpretable symbols emerge in the learned codebook of V3, showing its ability of abstracting concepts from information. To evaluate the interpretability of learned content representations quantitatively, we propose two metrics: the learned content codebook accuracy and standard deviation among styles. We first align codebook entries to the ground truth content labels by their distributions on content labels, and then calculate the accuracy of codebook entries’ distribution regarding their aligned labels. A well learned interpretable codebook should have entries concentrated on content labels they are aligned with, thus showing high accuracy. Also, as a good symbol is a symbol of consensus, on datasets with discrete style labels, we also quantify different styles’ discrepancy of codebook entries distribution on content labels using the standard deviation ($\sigma$) of confusion matrices between codebook entries and content labels, as shown in Table 8. For datasets with no discrete style labels, we report the accuracy of codebook entries in Table 9. Visualizations of confusion matrices between the codebook and ground truth content labels can be found in Appendix C.2.
From Table 8 and Table 9, we observe that V3 shows good codebook interpretability by representing content labels with consistent codebook entries, and the consistency is kept well among styles. Table 8 also shows that V3 shows good codebook interpretability with or without vocabulary redundancy, indicating V3 does not rely on the knowledge of the number of content classes to learn interpretable symbols.
Qualitatively, we perform content and style recombination by traversing all content codebook entries and decoding them with a fixed style representation. If the learned codebook has good interpretability, the decoding results of recombined content and style representations should show meaningful content changes, and retain consistent styles. Here, we focus on comparing the recombination results of V3 and baselines on SVHN, where we first encode ${\bm{z}}^{\mathrm{s}}$ from example fragments, and then recombine it with all $K$ content codebook entries. More experiment results on PhoneNums and InsNotes can be found in Appendix C.2 and our demo website.
Table 8: Quantitative results of codebook interpretability on datasets with discrete style labels. Values are reported in percentage.
     Method      PhoneNums      InsNotes
     $K$      Acc. $\uparrow$      $\sigma$downarrow      $K$      Acc. $\uparrow$      $\sigma$downarrow
     V3      10      89.2      0.6      12      99.8      0.1
     20      99.7      1.8      24      92.9      2.2
     40      99.9      4.5      48      90.2      4.5
     MINE-based      10      40.9      8.1      12      13.8      3.8
     20      25.6      9.8      24      29.4      8.9
     40      50.6      6.2      48      26.9      3.6
     Cycle loss      10      71.0      3.7      12      27.5      11.4
     20      89.6      4.3      24      28.5      11.9
     40      99.9      4.1      48      18.2      6.2
Table 9: Quantitative results of codebook interpretability on datasets without discrete style labels. Values are reported in percentage.
     Method      SVHN      Sprites      Libri100
     $K$      Acc. $\uparrow$      $K$      Acc. $\uparrow$      $K$      Acc. $\uparrow$
     V3      20      47.6      18      98.5      80      24.7
     MINE-based      20      26.0      18      38.3      80      10.8
     Cycle loss      20      20.1      18      82.0      80      10.5
From Figure 3, we observe that V3 generates images with clear content changes and consistent styles when recombining content and style representations. Although images generated by V3 may not cover all possible contents as many of them never appear in the training set in the given styles, the content are almost all recognizable digits and V3 can even “imagine” reasonably what a digit would look like in a new font and color. In contrast, the baselines generate images with either mixed content and style information, or very subtle changes in content that are hard to interpret. This comparison not only validates the interpretability of V3’s learned codebook resulted from successful content-style disentanglement, but also demonstrates the potential of V3 in style transfer and content editing tasks.
Figure 3: Comparison of generated images by recombining ${\bm{z}}^{\mathrm{s}}$ from given sources in SVHN and all ${\bm{z}}^{\mathrm{c}}$ in the learned codebook.
@@ 5 Limitation
We have identified several limitations in our V3 method that necessitate further investigation. First, while V3 achieves good disentanglement and symbolic interpretability, it is not flawless — samples of different contents (say images of “8" and “9") may be projected into the same latent code. Inspired by human learning, which effectively integrates both mode-1 and mode-2 cognitive processes, we aim to enhance V3 by incorporating certain feedback or reinforcement. This adaptation could also facilitate the application of V3 to more complex domains such as general image or video. Also, V3 focuses on learning the atomic content symbols that compose meanings of data samples, but the current approach does not learn the meanings of individual symbols or those emerging from specific symbol permutations. Additionally, V3 is currently optimized to disentangle content and style from data samples that include defined fragments. Extending this capability to unsegmented data of large vocabularies, such as continuous audio, represents a significant area for future development. Furthermore, V3 assumes that content elements do not overlap, which does not hold in cases of polyphonic music or mixed audio. Addressing this challenge will require a more sophisticated approach that considers the hierarchical nature of content.